OpenStudy (anonymous):

Write the equation of the line in point-slope form that passes through (5, -1) and has a slope of

OpenStudy (anonymous):

*of (2/3)

OpenStudy (anonymous):

use y=mx+b -1=2/3*5+b sfb -13/3 y=2/3 * x -13/3